if sum of two unit vector is a unit vector then magnitude of their difference  is

Grade:11

1 Answers

Priyanshu Gujjar
2 years ago

 

Let a,b be two unit vectors whose sum is also a unit vector c
 
a+b=c ...... (i) and a=b=c=1
 
Now, magnitude of sum of a and b is 
 
a+b=a2+b2+2(ab) 
 
1=1+1+2(ab) ..... From (i)
 
ab=21
 
Magnitude of their difference is given by
 
ab2=a2+b22a.b
 
              =1+1+2×21=3
 
ab=